# Music Arena: Live Evaluation for Text-to-Music

### v1

@@ 1 Introduction
Text-to-music (TTM) generation has advanced rapidly in recent years, with models demonstrating remarkable capabilities in creating high-fidelity music audio from text prompts (Dhariwal et al., 2020; Forsgren and Martiros, 2022; Agostinelli et al., 2023; Copet et al., 2023; Evans et al., 2025). This progress has highlighted two critical and intertwined challenges for the research community. Firstly, designing rigorous TTM evaluation protocols is essential for navigating tradeoffs in methodologies and training data, and also to track progress over time. Secondly, identifying a source of open and renewable human TTM preference data would help researchers to better align TTM systems with human intent (Cideron et al., 2024), and aid in the development of more reliable automatic evaluation metrics (Grötschla et al., 2025; Huang et al., 2025).
The current TTM landscape is unable to meet these challenges. Music is a human endeavor, and a rigorous evaluation metric should thus reflect human preferences. However, human preferences are difficult to capture as they may be influenced more by subjective assessments of creativity than objective, quantifiable phenomena. While numerous automatic evaluation metrics have been proposed (Kilgour et al., 2019; Gui et al., 2024; Huang et al., 2025), past work shows they correlate imperfectly with human preferences (Huang et al., 2025; Grötschla et al., 2025) and may not capture all key musical desiderata (Huang et al., 2025). Moreover, while some open preference datasets have been released (Liu et al., 2025a; Huang et al., 2025; Grötschla et al., 2025), these one time efforts are not renewable and will remain fixed even as new models emerge or human preferences drift. Commercial providers may have access to renewable sources of preferences through their platform’s proprietary usage data, but this data is typically not open.
@@
A new evaluation protocol, which we refer to here as live evaluation, has already helped navigate analogous challenges in other AI domains (Chiang et al., 2024; mrfakename et al., 2025; Jiang et al., 2024; Ebert, 2025; Chi et al., 2025). The key idea behind live evaluation is to align incentives by offering everyday users free access to generative AI systems in exchange for their preferences. On most live evaluation platforms, users first submit an input query, are presented with outputs from two different AI systems, and finally asked which of the two they prefer. These pairwise preferences are distilled into a global leaderboard ordered by Bradley-Terry coefficients (Bradley and Terry, 1952) or related scores like Elo (Elo, 1967). Compared to ad hoc human evaluation protocols, these scores are more consistent because they are calculated for all models from the same protocol and annotator distribution. Moreover, collecting preferences in this fashion is scalable because incentives are aligned, and the preference data reflects more realistic usage. Live evaluation was first proposed for language via Chatbot Arena (Chiang et al., 2024), and subsequently explored for text-to-speech (mrfakename et al., 2025), image and video generation (Jiang et al., 2024), 3D model generation (Ebert, 2025), and coding assistance (Chi et al., 2025). Here we propose to offer live evaluation for TTM.
@@
We also implement policies for Music Arena that aim to both increase platform trust and provide renewable access to preference data. Firstly, we anonymize private user information such as IP adddresses by salting and hashing, ensuring user privacy while also facilitating record linkage for longitudinal preference research. Secondly, our entire platform is open source [^1] Code available at https://github.com/gclef-cmu/music-arena aside from production configurations and secret keys. Finally, we propose to release data at regular (monthly) intervals, allowing the research community to access the latest data. By open sourcing our code and data, the entire lifecycle of our platform and leaderboard can be independently audited.
We release Music Arena to address key challenges in the landscape of TTM. By offering a consistent collection protocol, scalable alignment of incentives, and a real-world distribution of users, preferences collected via Music Arena may reflect a more rigorous evaluation protocol for TTM. Through a policy of consistent and privacy-preserving data releases, Music Arena aims to provide an open and renewable source of TTM preference data. Below we provide an overview of the Music Arena platform, including key features, policies, leaderboard structure, and ethical considerations.
@@ 2.1 Frontend
Our frontend is a web-based interface built with Gradio (Abid et al., 2019) and serves as the primary means of user interaction. Upon their first visit, users are presented with a consent page detailing the IRB-approved research protocol and data handling policies. Once consent is provided, the main “Arena” interface allows users to engage in one or more “Battles” (pairwise comparisons). To initiate a battle, a user submits a text prompt of their choosing. Next, the user is presented with two audio tracks generated by two different TTM systems, and may listen to them in whatever order and for however long they like. To mitigate potential biases that could arise from differing audio lengths, the interface supports variable-length outputs but conceals the specific duration of each track from the user. After listening, users specify their preference (prefer A, prefer B, tie, both bad). Finally, after voting, the frontend reveals the identities of the competing models, along with other information such as generation speed. A download link for the preferred track is provided as an incentive for casting a decisive vote, and users are given the option to provide additional natural language feedback.
@@ 2.3 Model Endpoints
To facilitate Music Arena, we aim to unify the heterogeneous type signatures of TTM systems. To this end, we program a model endpoint for numerous TTM systems (Copet et al., 2023; Evans et al., 2025; Liu et al., 2025b; Gong et al., 2025; DeepMind, 2025a, b; Forsgren and Martiros, 2022)—code that adapts underlying type signatures and dependencies into a common interface. For open weights models, endpoints manage synchronous calls to GPU resources and batching for increased throughput. For API-based commercial systems, endpoints adapt inputs from our unified API to HTTP requests to proprietary APIs. To manage the varied (and sometimes conflicting) software dependencies across systems, each endpoint is paired with a bespoke Docker container. Each container exposes a simple API endpoint with a common type signature, allowing the backend to interact with all systems in a uniform manner. This modular architecture facilitates decentralized development—providers can contribute new TTM systems to Music Arena without disturbing other parts of the platform.
@@ 3.1 LLM-based moderation and routing
The backend’s orchestration is powered by an LLM-based system that facilitates moderation of malicious input prompts, and routing of prompts on a unified interface (simple textbox) to relevant models. To moderate, the LLM is instructed to reject the user’s input prompt if it contains references to copyrighted musical material, culturally insensitive themes, or explicit themes, including profanity that would be atypical for the musical style (e.g., profanity okay for heavy metal, not okay for a nursery rhyme). For prompts that pass moderation, the LLM is instructed to extract structured information from the natural language input: the implied presence of vocals or lyrics (e.g., “folk song about a cat named Chamomile” implies lyrics), and explicitly-specified duration (e.g., “303030 second lo-fi beat”). The backend uses this structured representation to seamlessly route prompts to a subset of models that support the user’s query (e.g., many models do not support vocal or lyrics generation). At time of writing, Music Arena uses OpenAI’s GPT-4o (Hurst et al., 2024) for this component.
@@ 3.2 Detailed preferences via listening data and language feedback
Most live evaluation platforms for other AI domains collect simple pairwise preferences. Here we propose to additionally collect more detailed preference signals including fine-grained listening data and natural language feedback. As the user listens to each generated audio during a battle, our system stores their listening behavior including the amount of time spent listening to each clip, and the wall clock time at which they played or paused each clip. To ensure meaningful user engagement, the voting interface is only unlocked after a user listens to each track for a predefined minimum duration (444 seconds at the time of writing). After a user specifies their preference between four options—A is better, B is better, Tie, or Both are bad—they are encouraged to provide additional natural language feedback clarifying their rationale. We hope to use listening data to better understand and model user behavior, and the language feedback to offer richer insights into preferences and desiderata than binary preferences alone can provide. A full example of the detailed preference data we collect per battle appears in Appendix˜A.
@@ 3.3 Reference TTM implementations
A key feature of Music Arena is the development of a unified Docker-based framework for managing inference from TTM systems as outlined in Section˜2.3. In addition to supporting our core platform, we hope that this unified framework may benefit other research that requires comparing outputs from several TTM systems. At time of writing, we support the following open weights models: Meta’s MusicGen (Copet et al., 2023), Stability AI’s Stable Audio Open (Evans et al., 2025) and Stable Audio Open Small (Novack et al., 2025), SongGen (Liu et al., 2025b), ACE Studio’s ACE-Step (Gong et al., 2025), and Google DeepMind’s Magenta RealTime (DeepMind, 2025a). We also support API-based commercial models including Riffusion’s FUZZ models (1.0 & 1.1) (Team, 2025), Stability AI’s Stable Audio 2.0 (Evans et al., 2024), and Google DeepMind’s Lyria RealTime (DeepMind, 2025b). Due to resource limitations, not all of these models will be available for live evaluation at a given time, however they can always be accessed by researchers running our code using their own resources.
Systems in this collection exhibit substantial heterogeneity in type signatures. Three support generating output vocals (Liu et al., 2025b; Gong et al., 2025; Team, 2025) while others are instrumental only. Commercial systems like Riffusion FUZZ (Team, 2025) generate lyrics jointly with audio, while open weights systems like SongGen (Liu et al., 2025b) and ACE-Step (Gong et al., 2025) require explicit lyrics conditioning—here we use GPT-4o (Hurst et al., 2024) to generate lyrics for these systems from a user’s input prompt. In addition to considerations around vocals, there is a long tail of additional control signals across models, e.g., Stability AI models support explicit specification of output duration (Evans et al., 2025; Novack et al., 2025; Evans et al., 2024). This heterogeneity is particularly pronounced in music—other generative AI tasks like chat (Chiang et al., 2024) and image generation (Jiang et al., 2024) have more standardized type signatures. We design our unified framework to navigate this complex landscape, and aim to extend it in future work to support even broader music type signatures beyond TTM such as style transfer or symbolic music generation.
@@ 4 Key Policies
Along with the key features outlined previously, here we emphasize key policies of Music Arena designed to increase platform trust by promoting user privacy and platform transparency, and to provide renewable access to preference data.
@@ 4.1 Protecting user privacy while facilitating record linkage
The ability to perform record linkage across Music Arena sessions—identifying multiple battles from the same user—is critical for longitudinal preference analysis and detecting spam or malicious behavior. However, it is essential that we also protect user privacy by ensuring that personally identifying information is never exposed.
Following established recommendations for user privacy in research (Tene and Polonetsky, 2012), we implement a pseudonymization protocol using salted hashing (Kushida et al., 2012). When users interact with Music Arena, we transform linkable identifiers such as IP addresses by applying a server-side salt (a secret random string) followed by a one-way cryptographic hash. We only store these anonymized user identifiers, never the original identifiers. This approach provides strong privacy guarantees and protects against de-anonymization strategies such as rainbow table attacks. The resulting anonymized identifiers allow anyone to link battles from the same user across sessions for research purposes without ever exposing their private identifiers.
@@ 4.2 Maximizing platform transparency and data access
We are committed to making Music Arena as transparent as possible. To this end, all of our platform code is open source, aside from secret keys for private salting and API access. Additionally, we commit to a policy of rolling, comprehensive data releases. Unlike preferences from one-time data collection efforts (Liu et al., 2025a; Huang et al., 2025; Grötschla et al., 2025), we aim to publish Music Arena data at regular monthly intervals. This rolling approach is critical in the rapidly evolving field of generative AI, addressing key sources of distribution shift like the development of new TTM systems, and changes in user preferences over time. Moreover, we aim for our data releases to be comprehensive, [^2] Minor exceptions to this policy may apply, e.g., some model licenses prevent the release of generated audio. including anonymized user identifiers, generated audio, and detailed preferences. Through open source code and comprehensive data releases, we allow anyone in the research community to audit the entire lifecycle of our platform and evaluation results.
5 Leaderboard structure
We are releasing Music Arena alongside this paper. Accordingly, at time of writing, we do not have sufficient data to construct a public leaderboard. Here we provide an overview of our leaderboard plans, which we will release once a sufficient number of votes have been collected.
Public leaderboards for live evaluation platforms (Chiang et al., 2024; Jiang et al., 2024; Chi et al., 2025; mrfakename et al., 2025) tend to contain similar attributes: an overall “Arena Score” (usually derived from the Bradley-Terry model (Bradley and Terry, 1952)), the number of votes, the model provider, and the system license. There are a number of unique considerations in music that motivate presentation of additional attributes.
For LLM training, it can be broadly assumed that all providers are training on large-scale text data mined from the web. However, in music, we see significantly more diversity in training data across models. For example, some models are trained on licensed stock music (Copet et al., 2023) or publicly-available music under Creative Commons licenses (Evans et al., 2025), while some only specify the quantity (rather than the provenance) of their training data (Gong et al., 2025). These differences induce an uneven playing field for training data, affecting not only a system’s performance in pairwise comparisons but also its standing within broader legal and ethical conversations. Accordingly, we plan to include on our leaderboard a summary of available training data information for each model, including provenance and quantity.
In addition to training data, we plan to include generation speed on our leaderboard. Tools that facilitate the creation of music are often designed to have low latency interactions. Accordingly, TTM providers may make different tradeoffs between quality and speed, depending on if they are targeting more consumption-oriented (prefer quality) or creative-oriented (prefer speed) applications. Speed will be codified on our leaderboard by median real-time factor (RTF), where RTF is the ratio of seconds of music generated divided by seconds of wall clock time to generate. For example, a system that generates 303030 seconds of audio in 333 seconds has an RTF of 101010x. Measuring speed via RTF gracefully handles variable length outputs, ensuring that systems are not punished for taking more time to generate longer audio.
We will allow viewers to sort and filter the leaderboard by these additional characteristics, allowing viewers to make principled decisions about models based on attributes of keen importance to their specific music application goals. Moreover, we aim to further emphasize these music-specific tradeoffs through visualization: a  [^2] D scatter plot with speed on the X axis, Arena Score on the Y axis, and colors and shapes to emphasize training data and licensing information.
@@ 6 Ethical Considerations and Safeguards
Training data disclosure. Training data is at the center of legal and ethical conversations around the landscape of music generation (Barnett, 2023). We plan to surface available training data information on our public leaderboard, empowering the community to make more holistic assessments of TTM systems.
There remain ethical considerations for our work beyond these safeguards. Music Arena provides increased access to TTM systems for everyday users, which could have long term psychological effects. Our user distribution will likely be skewed to US users and AI enthusiasts, potentially promoting increased focus to the needs of those user populations by model providers. Music Arena currently focuses on text-to-music rather than other music generation tasks (e.g., style transfer, symbolic generation), which may shift focus to text-to-music and away from other tasks. Preferences on Music Arena may reflect quality more than diversity, which may ultimately reduce the musical diversity across TTM systems.
Music Arena also inherits many ethical and societal considerations from music generation more broadly. Music generation may change the economic landscape of music labor, accelerate the commodification of music, or contribute to the homogenization of music cultures. Overall, we believe the benefits of Music Arena (more rigorous and transparent evaluation, open availability of preference data) outweigh the risks. Additionally, a strength of live evaluation is that it can be stopped at any time—we will monitor broader ethical considerations as the platform develops.

### v2

@@ Abstract
Preference data is available at: https://huggingface.co/music-arena
@@ 1 Introduction
Text-to-music (TTM) generation has advanced rapidly in recent years, with models demonstrating remarkable capabilities in creating high-fidelity music audio from text prompts [1, 2, 3, 4, 5]. This progress has highlighted two critical and intertwined challenges for the research community. Firstly, designing rigorous TTM evaluation protocols is essential for navigating tradeoffs in methodologies and training data, and also to track progress over time. Secondly, identifying a source of open and renewable human TTM preference data would help researchers to better align TTM systems with human intent [6], and aid in the development of more reliable automatic evaluation metrics [7, 8].
The current TTM landscape is unable to meet these challenges. Music is a human endeavor, and a rigorous evaluation metric should thus reflect human preferences. However, human preferences are difficult to capture as they may be influenced more by subjective assessments of creativity than objective, quantifiable phenomena. While numerous automatic evaluation metrics have been proposed [9, 10, 8], past work shows they correlate imperfectly with human preferences [8, 7] and may not capture all key musical desiderata [8]. Moreover, while some open preference datasets have been released [11, 8, 7], these one time efforts are not renewable and will remain fixed even as new models emerge or human preferences drift. Commercial providers may have access to renewable sources of preferences through their platform’s proprietary usage data, but this data is typically not open.
@@
A new evaluation protocol, which we refer to here as live evaluation, has already helped navigate analogous challenges in other AI domains [12, 13, 14, 15, 16]. The key idea behind live evaluation is to align incentives by offering everyday users free access to generative AI systems in exchange for their preferences. On most live evaluation platforms, users first submit an input query, are presented with outputs from two different AI systems, and finally asked which of the two they prefer. These pairwise preferences are distilled into a global leaderboard ordered by Bradley-Terry coefficients [17] or related scores like Elo [18]. Compared to ad hoc human evaluation protocols, these scores are more consistent because they are calculated for all models from the same protocol and annotator distribution. Moreover, collecting preferences in this fashion is scalable because incentives are aligned, and the preference data reflects more realistic usage. Live evaluation was first proposed for language via Chatbot Arena [12], and subsequently explored for text-to-speech [13], image and video generation [14], 3D model generation [15], and coding assistance [16]. Here we propose to offer live evaluation for TTM.
@@
We also implement policies for Music Arena that aim to both increase platform trust and provide renewable access to preference data. Firstly, we anonymize private user information such as IP addresses by salting and hashing, ensuring user privacy while also facilitating record linkage for longitudinal preference research. Secondly, our entire platform is open source [^1] Code is available at https://github.com/gclef-cmu/music-arena aside from production configurations and secret keys. Finally, we propose to release data at regular (monthly) intervals, allowing the research community to access the latest data. By open sourcing our code and data, the entire lifecycle of our platform and leaderboard can be independently audited.
We release Music Arena to address key challenges in the landscape of TTM (rigorous evaluation, renewable preference data). Here, we present an overview of the Music Arena platform alongside an analysis of the preferences collected between July 28 and Aug 31, 2025, including 1,420 user-initiated battles from 373 unique users yielding 1,051 votes. From these votes, we build a preliminary Music Arena leaderboard (Appendix˜A), ranking contemporary TTM models via live evaluation.
@@ 2.1 Frontend
Our frontend is a web-based interface built with Gradio [19] and serves as the primary means of user interaction. Upon their first visit, users are presented with a consent page detailing the IRB-approved research protocol and data handling policies. Once consent is provided, the main “Arena” interface allows users to engage in one or more “Battles” (pairwise comparisons). To initiate a battle, a user submits a text prompt of their choosing. Next, the user is presented with two audio tracks generated by two different TTM systems, and may listen to them in whatever order and for however long they like. To mitigate potential biases that could arise from differing audio lengths, the interface supports variable-length outputs but conceals the specific duration of each track from the user. After listening, users specify their preference (prefer A, prefer B, tie, both bad). Finally, after voting, the frontend reveals the identities of the competing models, along with other information such as generation speed. A download link for the preferred track is provided as an incentive for casting a decisive vote, and users are given the option to provide additional natural language feedback.
@@ 2.3 Model Endpoints
To facilitate Music Arena, we aim to unify the heterogeneous type signatures of TTM systems. To this end, we implement a model endpoint for numerous TTM systems [4, 5, 20, 21, 22, 23, 2]—code that adapts underlying type signatures and dependencies into a common interface. For open weights models, endpoints manage synchronous calls to GPU resources and batching for increased throughput. For API-based commercial systems, endpoints adapt inputs from our unified API to HTTP requests to proprietary APIs. To manage the varied (and sometimes conflicting) software dependencies across systems, each endpoint is paired with a bespoke Docker container. Each container exposes a simple API endpoint with a common type signature, allowing the backend to interact with all systems in a uniform manner. This modular architecture facilitates decentralized development—providers can contribute new TTM systems to Music Arena without disturbing other parts of the platform.
@@ 3.1 LLM-based moderation and routing
The backend’s orchestration is powered by an LLM-based system that facilitates moderation of malicious input prompts, and .0routing of prompts on a unified interface (simple textbox) to relevant models. To moderate, the LLM is instructed to reject the user’s input prompt if it contains references to copyrighted musical material, culturally insensitive themes, or explicit themes, including profanity that would be atypical for the musical style (e.g., profanity okay for heavy metal, not okay for a nursery rhyme). For prompts that pass moderation, the LLM is instructed to extract structured information from the natural language input: the implied presence of vocals or lyrics (e.g., “folk song about a cat named Chamomile” implies lyrics), and explicitly-specified duration (e.g., “$30$ second lo-fi beat”). The backend uses this structured representation to seamlessly route prompts to a subset of models that support the user’s query (e.g., many models do not support vocal or lyrics generation). At time of writing, Music Arena uses OpenAI’s GPT-4o [24] for this component.
@@ 3.2 Detailed preferences via listening data and language feedback
Most live evaluation platforms for other AI domains collect simple pairwise preferences. Here we propose to additionally collect more detailed preference signals including fine-grained listening data and natural language feedback. As the user listens to each generated audio during a battle, our system stores their listening behavior including the amount of time spent listening to each clip, and the wall clock time at which they played or paused each clip. To ensure meaningful user engagement, the voting interface is only unlocked after a user listens to each track for a predefined minimum duration ($4$ seconds at the time of writing). After a user specifies their preference between four options—A is better, B is better, Tie, or Both are bad—they are encouraged to provide additional natural language feedback, clarifying their rationale. We use listening data to better understand and model user behavior, and the language feedback to offer richer insights into preferences and desiderata than binary preferences alone can provide. An example of our detailed preference data per battle and an analysis of the initial data release appear in  Appendix˜C and  Appendix˜B, respectively.
@@ 3.3 Reference TTM implementations
A key feature of Music Arena is the development of a unified Docker-based framework for managing inference from TTM systems as outlined in Section˜2.3. In addition to supporting our core platform, we hope that this unified framework may benefit other research that requires comparing outputs from several TTM systems. At time of writing, we support the following open weights models: Meta’s MusicGen [4], Stability AI’s Stable Audio Open [5] and Stable Audio Open Small [25], SongGen [20], ACE Studio’s ACE-Step [21], and Google DeepMind’s Magenta RealTime [22]. We also support API-based commercial models including Producer.ai’s FUZZ models (1.0 & 1.1) [26], Stability AI’s Stable Audio 2.0 [27], and Google DeepMind’s Lyria RealTime [23]. Due to resource limitations, not all of these models will be available for live evaluation at a given time, however they can always be accessed by researchers running our code using their own resources.
Systems in this collection exhibit substantial heterogeneity in type signatures. Three support generating output vocals [20, 21, 26] while others are instrumental only. Commercial systems like Riffusion FUZZ [26] generate lyrics jointly with audio, while open weights systems like SongGen [20] and ACE-Step [21] require explicit lyrics conditioning— we use GPT-4o [24] to generate lyrics for these systems from a user’s input prompt. In addition to considerations around vocals, there is a long tail of additional control signals across models, e.g., Stability AI models support explicit specification of output duration [5, 25, 27]. This heterogeneity is more pronounced in music, unlike the standardized signatures of chat [12] and image generation [14]. We design our unified framework to navigate this complex landscape, and aim to extend it in future work to support even broader music type signatures beyond TTM such as style transfer or symbolic music generation.
@@ 4 Key Policies
Along with the key features outlined previously, here we emphasize key policies of Music Arena designed to increase platform trust by surfacing considerations beyond user preferences, promoting user privacy and platform transparency, and providing renewable access to preference data.
4.1 Surfacing model considerations beyond Arena Score
The primary metric of interest across existing live evaluation platforms in other domains is an “Arena Score”, derived from pairwise comparisons. In music, holistic comparisons may require considerations of factors beyond preferences. Accordingly, our leaderboard (Appendix˜A) also surfaces information on training data and generation speed. Including information on training data acknowledges the legal, ethical, and quality implications of the “uneven playing field” for training data across model providers. Reporting generation speed (measured by median RTF) recognizes that some models may trade off quality for speed, e.g., to facilitate low-latency creative workflows.
@@ 4.2 Protecting user privacy while facilitating record linkage
The ability to perform record linkage across Music Arena sessions—identifying multiple battles from the same user—is critical for longitudinal preference analysis and detecting spam or malicious behavior. However, it is essential that we also protect user privacy by ensuring that personally identifying information is never exposed. Following established recommendations for user privacy in research [28], we implement a pseudonymization protocol using salted hashing [29]. When users interact with Music Arena, we transform linkable identifiers such as IP addresses by applying a server-side salt (a secret random string) followed by a one-way cryptographic hash. We only store these anonymized user identifiers, never the original identifiers. This approach provides strong privacy guarantees and protects against de-anonymization strategies such as rainbow table attacks. The resulting anonymized identifiers allow anyone to link battles from the same user across sessions for research purposes without ever exposing their private identifiers.
@@ 4.3 Maximizing platform transparency and data access
We are committed to making Music Arena as transparent as possible. To this end, all of our platform code is open source, aside from secret keys for private salting and API access. Additionally, we commit to a policy of rolling, comprehensive data releases. Unlike preferences from one-time data collection efforts [11, 8, 7], we aim to publish Music Arena data at regular monthly intervals. This rolling approach is critical in the rapidly evolving field of generative AI, addressing key sources of distribution shift like the development of new TTM systems, and changes in user preferences over time. Moreover, we aim for our data releases to be comprehensive, [^2] Minor exceptions to this policy may apply, e.g., some model licenses prevent the release of generated audio. including anonymized user identifiers, generated audio, and detailed preferences. This combination of open code and data allows the research community to audit our platform’s entire lifecycle and evaluation results.
@@ 5 Ethical Considerations and Safeguards
There remain ethical considerations for our work beyond these safeguards. Music Arena provides increased access to TTM systems for everyday users, which could have long term psychological effects. Our user distribution will likely be skewed to US users and AI enthusiasts, potentially promoting increased focus to the needs of those user populations by model providers. Music Arena’s current focus on text-to-music may inadvertently steer the research community’s attention away from other important tasks like style transfer or symbolic generation. Music Arena also inherits many ethical and societal considerations from music generation more broadly. Music generation may change the economic landscape of music labor, accelerate the commodification of music, or contribute to the homogenization of music cultures. Overall, we believe the benefits of Music Arena (more rigorous and transparent evaluation, open availability of preference data) outweigh the risks.